prove that every non trival subgroup of a cyclic group has finite index hence prove that (Q,+) is non cyclic


1
Expert's answer
2021-05-07T11:53:41-0400

Solution:

Given, G is a cyclic group.

Therefore, G=<a>.

And H is subgroup of G.

H=<"a^i" >.

Index,

G/H={"a^j+<a^i>" }

If j>i then by division algorithm, j=ir+s

Then "a^j+<a^i>=a^s+<a^i>"

So, G/H={"a^s+<a^i>, 0<=s<i" }

O(G/H)= finite =index of H.

For( Q,+), choose H=<1/2>

Clearly o(G/H)= inifinte. 

So, G=(Q,+) is not cyclic